# Explore Synergistic Interaction Across Frames for Interactive Video Object Segmentation

### v1

@@ VI Conclusion
iVOS is a human-computer interaction task that requires the quality of user experience, but previous methods are limited to single input mode and slow running speed. Thus, we propose SIAF, a novel multi-frame and multi-object framework that allows the users to interact with multiple frames at once, which is a pioneer of multi-frame interaction. In detail, SIAF achieves new SOTA performance on DAVIS 2017 (89.6% $\mathcal{J\&F}@60$) when across-frame scribbles are available. Meanwhile, our R50-SIAF is more than $\bf{3\times}$ faster than the SOTA competitor under challenging multi-object scenarios. Additionally, we introduced an interactive graphical user interface (GUI) that supports multi-frame interaction. We hope our SIAF can serve as a solid multi-frame & multi-object baseline in iVOS and will promote the use of iVOS in real-world applications.
Broader Impact and Limitations. 1) The efficacy of iVOS methods is largely contingent on input scribble quality. In cases where the input scribbles are rough, segmentation performance may suffer. The sensitivity of scribble inputs analysis can be found in supplementary materials. 2) SIAF is meticulously engineered to concurrently process multiple frames and objects, presenting an optimal solution for forthcoming tracking requirements aimed at minimizing time consumption.

### v3

@@ VI Conclusion
iVOS is a human-computer interaction task that requires the quality of user experience, but previous methods are limited to single input mode and slow running speed. Thus, we propose IDPro, a novel multi-frame and multi-object framework that allows the users to interact with multiple frames at once, which is a pioneer of multi-frame interaction. In detail, IDPro achieves new SOTA performance on DAVIS 2017 dataset (89.6% $\mathcal{J\&F}@60$) when across-frame scribbles are available. Meanwhile, our R50-IDPro is more than $\bf{3\times}$ faster than the SOTA competitor under challenging multi-object scenarios. Concretely, we proposed an across-frame interaction module that facilitates the transfer of scribble-based information between interactive frames. We develop a concurrent propagation module that can simultaneously process multiple objects, and design a re-propagation strategy to exploit the interactive information from multiple rounds. Additionally, we introduced an interactive graphical user interface (GUI) that supports multi-frame interaction. We hope our IDPro can serve as a solid multi-frame and multi-object baseline in iVOS and will promote the use of iVOS in real-world applications.
Broader Impact and Limitations. 1) The efficacy of iVOS methods is largely contingent on input scribble quality. In cases where the input scribbles are rough, segmentation performance may suffer. 2) IDPro is meticulously engineered to concurrently process multiple frames and objects, presenting an optimal solution for forthcoming tracking requirements aimed at minimizing time consumption.
Acknowledgments
This work was supported by the National Key Research & Development Project of China (2021ZD0110700), the National Natural Science Foundation of China (62337001), the Fundamental Research Funds for the Central Universities(226202400058) and the Fundamental Research Funds for the Central Universities (No. 226-2022-00051).